# Task Aware Dreamer for Task Generalization in Reinforcement Learning

### v4

@@ 4.2 Optimization
Based on the above analyses, we now introduce the training of TAD in detail. Following previous world models (Hafner et al. 2019a), we adopt an alternating training approach between the reward-informed world models and the policy.
To balance different tasks, when collecting data, TAD utilizes $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$ to store trajectories sampled from $\{\mathcal{M}_{m}\}_{m=1}^{M}$, respectively. Then, TAD samples data from each replay buffer and trains the reward-informed world models via the optimization objective following Eq. (LABEL:eq_elbo) as
$\begin{split}&L_{\text{TAD}}=\sum_{i=1}^{M}\mathbb{E}_{q}\left[\sum_{t=1}^{T}\ln p_{\theta}(o_{t}^{i}|h_{t}^{i},s_{t}^{i})+\sum_{t=1}^{T}\ln p_{\theta}(r_{t}^{i}|h_{t}^{i},s_{t}^{i})\right.\\ &-\left.\sum_{t=1}^{T}D_{\text{KL}}(q(s_{t}^{i}|h_{t}^{i},o_{t}^{i})\|p_{\theta}(s_{t}^{i}|h_{t}^{i}))+L_{\text{task}}\right].\end{split}$ (4)
In Eq. (LABEL:eq_loss), the first three items are similar to Dreamer for reconstructing observations, predicting rewards, and inferring states. Besides them, TAD includes an additional task term $L_{\text{task}}$ for predicting different tasks and learning task-aware embedding. In detail, we provide two alternatives: cross-entropy and self-contrastive, which are introduced below (We theoretically demonstrate their effectiveness in Sec. 4.3).
@@
The last term in Eq. LABEL:eq_elbo indicates that we need to maximize the log probability of the task context to distinguish different tasks via historical information. Thus TAD-CE directly maximizes the log probability over different tasks represented by one-hot vectors following previous works (Yang et al. 2020) and $L_{\text{task}}$ is set as below.
@@
Besides directly maximizing the log probability, we further propose TAD-SC, borrowing the idea of supervised contractive learning (Khosla et al. 2020). In detail, TAD-SC keeps the task embeddings of the same task closer and the task embeddings of different tasks far apart, which benefits reward-informed world models to better distinguish different tasks. Formally, we assume that the task model maps all sampled data as $\{m_{j}\}_{j=1}^{M\times T}$ and set $L_{\text{task}}$ as
@@
here $A(j)$ is the set of indices that are sampled with the same task of $m_{j}$, and $\tau$ is the temperature parameter, which is set as 0.1 following previous works (Khosla et al. 2020).
In terms of training the actor-critic that are parameterized neural networks, we extend the actor-critic learning in Dreamer to the task distribution. We first sample a series of states from the replay buffer and start from them to imagine trajectories via our reward-informed world models (results are in Fig. 3), which can capture invariant features and are beneficial for the agent to gain better generalization. After obtaining imagined trajectories, the actor-critic is optimized via maximizing the $\lambda$-return (Schulman et al. 2015) and regressing the TD targets (Sutton and Barto 2018), respectively.
@@
0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize parameters of world models, the policy, and the critic.
[…]
2:     for update $\text{update step}=1,2,...,U$ do
@@ 4.3 Theoretical Analyses
Below we introduce $3$ types of widely used policy hypotheses and show that $\Pi_{3}$, used in TAD, is necessary for handling task generalization.
@@
    Markovian policy set $\Pi_{1}$ (Sutton and Barto 2018; Yarats et al. 2021), i.e., $\Pi_{1}=\{\pi|\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})\}$, here $\Delta(\mathcal{A})$ represents a distribution over $\mathcal{A}$, which is widely used and optimal for the single-task setting;
@@
    $\mathcal{S}$-$\mathcal{A}$ memorized policy set $\Pi_{2}$ (Hafner et al. 2019a, 2020; Lee et al. 2020), i.e., $\Pi_{2}=\{\pi|\pi:\mathcal{H}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{H}=\cup_{t=1}^{\infty}\mathcal{H}_{t},\mathcal{H}_{t}=(\mathcal{S}\times\mathcal{A})^{t-1}\times\mathcal{S}$;
@@
    $\mathcal{S}$-$\mathcal{A}$$-\mathcal{R}$ memorized policy set $\Pi_{3}$ (Zintgraf et al. 2019; Rimon et al. 2024), i.e., $\Pi_{3}=\{\pi|\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$.
As illustrated in Fig. 1, naturally $\Pi_{1}\subseteq\Pi_{2}\subseteq\Pi_{3}$. Now we will analyze their expressive ability under task distribution $\mathcal{T}$. Denote $J_{\mathcal{T}}^{*}\triangleq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi}J_{\mathcal{M}}(\pi)\right]$ as the optimal return under $\mathcal{T}$, and $J_{\mathcal{T}}^{i}\triangleq\max_{\pi\in\Pi_{i}}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\right]$ as the optimal return for $\Pi_{i}(i=1,2,3)$ under $\mathcal{T}$. Our first result shows that, although $\Pi_{1}\subseteq\Pi_{2}$, they own the same expressive ability, i.e., $J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}$, and are both sub-optimal:
Theorem 2 (Sub-Optimality of $\Pi_{1},\Pi_{2}$. Proof in Appendix A.3).
We set ¯$\bar{\mathcal{M}}=(\mathcal{S},\mathcal{A},\mathcal{P},\bar{\mathcal{R}},\gamma)$, here ¯$\bar{\mathcal{R}}=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\mathcal{R}_{\mathcal{M}}]$. For $\forall\pi\in\Pi_{2}$, we have $\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\bar{\mathcal{M}}}(\pi)$ and further $J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}\leq J_{\mathcal{T}}^{*}$.
Theorem 2 reveals that the cumulative returns of policies in $\Pi_{1}$ and $\Pi_{2}$ are the same as their returns in the “average” MDP $\bar{\mathcal{M}}$, where the reward function is the average of reward functions in different tasks. Also, as $\Pi_{1}$ and $\Pi_{2}$ only choose actions via current state or historical state-action pairs, they cannot distinguish different tasks and are both sub-optimal.
To quantitatively analyze the characteristic of $\mathcal{T}$ and the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$, we propose a novel metric Task Distribution Relevance (TDR) of the distribution $\mathcal{T}$ as
Definition 1 (Task Distribution Relevance).
For any task distribution $\mathcal{T}$ and state $s$, the Task Distribution Relevance of $\mathcal{T}$ and $s$ is defined as
$\begin{split}D_{\text{TDR}}(\mathcal{T},s)&=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\max_{a}Q_{\mathcal{M}}^{*}(s,a)]\\ &-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[Q_{\mathcal{M}}^{*}(s,a)\right].\end{split}$ (7)
Intuitively, TDR describes the relevance of $\mathcal{T}$ via optimal $Q$ functions, which determine the distribution of optimal actions in corresponding tasks. Based on TDR, we can bound the gap:
Theorem 3 (Proof in Appendix A.4).
@@
$\begin{split}J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}[D_{\text{TDR}}(\mathcal{T},s)]\end{split}$ (8)
@@
Theorem 3 demonstrates that the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$ is related to the TDR of $\mathcal{T}$. When considering $\mathcal{T}$ with high TDR, i.e., the optimal $Q$ values in different tasks differ greatly, the performance of $\Pi_{1}$ and $\Pi_{2}$ will be extremely poor since they cannot differentiate different tasks and their expressive abilities are significantly limited. This conclusion is further verified empirically in the experimental section. Moreover, we can show that $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ might be arbitrarily small when $J_{\mathcal{T}}^{3}$ might be arbitrarily close to $J_{\mathcal{T}}^{*}$ (More details are in Appendix A.5), which demonstrates that $\Pi_{3}$ owns stronger expressive ability. Consequently, it is necessary to utilize $\Pi_{3}$ to distinguish different tasks for enhancing expression ability and generalization over the task distribution.
@@
Now we will show that, although $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ is a simple term, optimizing it can be effective for reaching the generalizable agent in the task distribution.
@@
As the input space of $\Pi_{3}$ is $\mathcal{L}$, i.e., all partial trajectories, our major result analyzes the relation between the policy $\pi\in\Pi_{3}$ and the task posterior $p(\mathcal{M}|l),l\in\mathcal{L}$. As $h_{t}$ encodes all historical information, $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ can be rewritten as $p(\mathcal{M}|l),l\in\mathcal{L}$.
Theorem 4 (Informally, detailed analyses and proof are in Appendix A.6).
For any policy $\pi\in\Pi_{3}$, we have
@@
$\begin{split}&J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\ -\left.\int_{a,\mathcal{M}}\pi(a|l)p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)dad\mathcal{M}\right]dl,\end{split}$ (9)
@@
Consequently, maximizing $p(\mathcal{M}|l)$, i.e., making the distribution of $p(\mathcal{M}|l)$ to be closer to some Dirac distribution, can significantly reduce the right part of Eq. (LABEL:eq_pi3), thus is effective for improving the generalization ability of $\pi$. More details and discussion are also provided in Appendix A.6.

### v5

@@ 4.2 Optimization
Based on the above analyses, we now describe the training of TAD in detail. Following prior world model approaches [5], TAD adopts an alternating training scheme between the RIWM and the policy.
To ensure balanced learning across tasks, TAD maintains $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$ to store trajectories collected from the corresponding task $\{\mathcal{M}_{m}\}_{m=1}^{M}$. During training, TAD samples data from each replay buffer and updates the RIWM by optimizing the objective in Eq. (3) as:
$\begin{split}&L_{\text{TAD}}=\sum_{i=1}^{M}\mathbb{E}_{q}\left[\sum_{t=1}^{T}\ln p_{\theta}(o_{t}^{i}|h_{t}^{i},s_{t}^{i})+\sum_{t=1}^{T}\ln p_{\theta}(r_{t}^{i}|h_{t}^{i},s_{t}^{i})\right.-\left.\sum_{t=1}^{T}D_{\text{KL}}(q(s_{t}^{i}|h_{t}^{i},o_{t}^{i})\|p_{\theta}(s_{t}^{i}|h_{t}^{i}))+L_{\text{task}}\right].\end{split}$ (5)
In Eq. (LABEL:eq_loss), the first three terms are similar to Dreamer for reconstructing observations, predicting rewards, and inferring states. In addition, TAD introduces a novel task context term $L_{\text{task}}$ to distinguish tasks and learn task-aware embeddings. Specifically, we propose two alternative formulations for this term: cross-entropy and self-contrastive, which are described in detail below. A theoretical justification for their effectiveness is provided in Sec. 4.3.
@@
The final term in Eq. (3) emphasizes the need to maximize the log-probability of the task context, enabling the model to distinguish tasks based on historical information. Following prior work [23], TAD-CE directly maximizes this log-probability over different tasks represented by one-hot vectors. The task loss $L_{\text{task}}$ is thus defined as:
@@
In addition to directly maximizing the log-probability, we further propose TAD-SC, inspired by supervised contrastive learning [43]. Specifically, TAD-SC encourages task embeddings from the same task to be close while pushing embeddings from different tasks apart, enabling RIWM to distinguish tasks more effectively. Formally, we assume that the task model maps all sampled data as $\{m_{j}\}_{j=1}^{M\times T}$ and define $L_{\text{task}}$ as:
@@
where $A(j)$ is the set of data indices that are collected from task $m_{j}$, and $\tau$ is the temperature parameter, which is set as 0.1 following previous works [43].
For training the actor-critic networks, we extend the actor-critic learning in Dreamer to the task distribution. Specifically, we first sample a set of states from the replay buffers and use our RIWM to imagine trajectories starting from these states (illustrated in Fig. 2). These imagined trajectories can capture task-specific features and are thus beneficial for the agent to gain better generalization. Once the imagined trajectories are obtained, the actor and critic are optimized by maximizing the $\lambda$-return [44] and regressing the TD targets [45], respectively.
@@
0:$M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, NN test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize parameters of world models, the policy, and the critic.
1:for $\text{iteration step}=1,2,$... do
2:  for $\text{update step}=1,2,$... do
@@ 4.3 Theoretical analyses
Below, we introduce $3$ types of widely used policy hypothesis sets and show that $\Pi_{3}$, used in TAD, is necessary for handling task generalization.
@@
    Markovian policy set $\Pi_{1}$ [45, 46], i.e., $\Pi_{1}=\{\pi|\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})\}$, where $\Delta(\mathcal{A})$ represents a distribution over $\mathcal{A}$, which is widely used and optimal for the single-task setting;
@@
    $\mathcal{S}$-$\mathcal{A}$ memorized policy set $\Pi_{2}$ [5, 32, 22], i.e., $\Pi_{2}=\{\pi|\pi:\mathcal{H}\rightarrow\Delta(\mathcal{A})\}$, where $\mathcal{H}=\cup_{t=1}^{\infty}\mathcal{H}_{t},\mathcal{H}_{t}=(\mathcal{S}\times\mathcal{A})^{t-1}\times\mathcal{S}$;
@@
    $\mathcal{S}$-$\mathcal{A}$-$\mathcal{R}$ memorized policy set $\Pi_{3}$ [31, 7], i.e., $\Pi_{3}=\{\pi|\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A})\}$, where $\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$.
As illustrated in Fig. 1a, these hypothesis sets naturally satisfy $\Pi_{1}\subseteq\Pi_{2}\subseteq\Pi_{3}$. We then analyze their expressive ability under task distribution $\mathcal{T}$. Let $J_{\mathcal{T}}^{*}\triangleq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi}J_{\mathcal{M}}(\pi)\right]$ denote the optimal return under $\mathcal{T}$, and $J_{\mathcal{T}}^{i}\triangleq\max_{\pi\in\Pi_{i}}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\right]$ the optimal return achievable by policies in $\Pi_{i}(i=1,2,3)$ under $\mathcal{T}$. Our first result shows that, although $\Pi_{1}\subseteq\Pi_{2}$, they own the same expressive ability, i.e., $J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}$, and are both sub-optimal:
Theorem 4.2 (Sub-Optimality of $\Pi_{1},\Pi_{2}$. Proof in Appendix A.3).
We set $\bar{\mathcal{M}}=(\mathcal{S},\mathcal{A},\mathcal{P},\bar{\mathcal{R}},\gamma)$, here $\bar{\mathcal{R}}=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\mathcal{R}_{\mathcal{M}}]$. For $\forall\pi\in\Pi_{2}$, we have $\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\bar{\mathcal{M}}}(\pi)$ and further $J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}\leq J_{\mathcal{T}}^{*}$.
Theorem 4.2 reveals that the cumulative returns of policies in $\Pi_{1}$ and $\Pi_{2}$ are equivalent as their returns in the “average” MDP $\bar{\mathcal{M}}$, where the reward function is the average of reward functions in different tasks. Moreover, since policies in $\Pi_{1}$ and $\Pi_{2}$ only select actions based on current state or historical state-action pairs, they are unable to distinguish different tasks and are thus both sub-optimal.
To quantitatively analyze the relationship between $\mathcal{T}$ and the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$, we propose a novel metric named Task Distribution Relevance (TDR), which characterize the distribution $\mathcal{T}$ as:
Definition 4.3 (Task Distribution Relevance).
For any task distribution $\mathcal{T}$ and state $s$, the Task Distribution Relevance of $\mathcal{T}$ and $s$ is defined as:
$\begin{split}D_{\text{TDR}}(\mathcal{T},s)&=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\max_{a}Q_{\mathcal{M}}^{*}(s,a)]-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[Q_{\mathcal{M}}^{*}(s,a)\right].\end{split}$ (8)
Intuitively, TDR quantifies the relevance of tasks in $\mathcal{T}$ via optimal $Q$ functions, which determine the choice of optimal actions in corresponding tasks. Based on TDR, we can derive a bound on the performance gap:
Theorem 4.4 (Regret of Markovian Policies).
@@
$\begin{split}J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}[D_{\text{TDR}}(\mathcal{T},s)].\end{split}$ (9)
@@
Theorem 4.4 demonstrates that the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$ is closely related to the TDR of $\mathcal{T}$. For task distributions $\mathcal{T}$ with high TDR, i.e., the optimal $Q$ values vary greatly across tasks, the performance of $\Pi_{1}$ and $\Pi_{2}$ can be extremely poor since they cannot differentiate between tasks and their expressive abilities are significantly limited. This conclusion is further confirmed empirically in the experimental section. Moreover, we show that $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ can be arbitrarily small when $J_{\mathcal{T}}^{3}$ can be arbitrarily close to $J_{\mathcal{T}}^{*}$ (see Appendix A.5 for details), demonstrating that $\Pi_{3}$ possesses stronger expressive ability. Consequently, it is necessary to utilize policies in $\Pi_{3}$ to distinguish different tasks and thus improve generalization across the task distribution.
@@
Now we will show that, although $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ is a simple term, optimizing it can be effective for obtaining a generalizable agent in the task distribution setting.
@@
As the input space of $\Pi_{3}$ is $\mathcal{L}$, i.e., all historical trajectories, our main result analyzes the relationship between a policy $\pi\in\Pi_{3}$ and the task posterior $p(\mathcal{M}|l),l\in\mathcal{L}$. As $h_{t}$ encodes all historical information, $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ can equivalently be expressed as $p(\mathcal{M}|l),l\in\mathcal{L}$.
Theorem 4.5 (Informal; detailed analyses and proof in Appendix A.6).
For $\forall\pi\in\Pi_{3}$, we have
@@
$\begin{split}J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]=&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.-\left.\int_{a,\mathcal{M}}\pi(a|l)p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)dad\mathcal{M}\right]dl,\end{split}$ (10)
where $p(l)$ is a distribution of $\mathcal{L}$ related to $\mathcal{T},\pi$ and $p(\mathcal{M}|l)$ is the task posterior related to $\pi$.
@@
Consequently, maximizing $p(\mathcal{M}|l)$, i.e., making the distribution of $p(\mathcal{M}|l)$ closer to a certain Dirac distribution, can significantly reduce the right part of Eq. (10), and thus is effective for improving the generalization ability of $\pi$. Additional details and discussion are provided in Appendix A.6.
5 Experiments
We now present empirical results to answer the following questions:
  • How well does TAD generalize across tasks when handling image-based and state-based observations? (Sec. 5.3)
  • Can TAD be extended to more general settings, such as dynamics generalization? (Sec. 5.4)
5.1 Experimental setup
Image-based control.
To verify our analyses of TDR in Sec. 4.3, we consider several task combinations in the DeepMind Control suite (DMC) [8]: (1) Cartpole-balance&balance_\_sparse, which shares the same optimal actions with TDR=0=0; (2) Walker-stand&walk&prostrat&flip, and Cheetah-run&run_\_back&flip&flip_\_flip_\_back, which are widely used in multi-task unsupervised RL [39, 47] and include task combinations with non-zero TDR. For example, Cheetah-run and Cheetah-flip hope a two-legged robot to move forward by running and flip around the torso, respectively (see Fig. 2), yielding distinct optimal Q-functions and thus leading to huge TDR. Additional details about these task combinations are provided in Appendix C.1.
To evaluate generalization with image-based observations, we extend classic tasks in DMC and design three task distributions: (1) Cheetah_\_speed($\alpha,\ \beta$), which extends Cheetah-run and hopes the agent to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$; (2) Pendulum_\_angle($\alpha,\ \beta$), extending Pendulum-swingup to keep pendulum’s pole within the target angle interval $(\arccos\alpha,\ \arccos\beta)$; and (3) Walker_\_speed($\alpha,\ \beta$), which is based on Walker-run and requires the planar walker to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$. For each task distribution, we sample $4$ training tasks and $2$ additional test tasks. More details are provided in Appendix C.2.
State-based control.
To demonstrate the scalability of TAD, we also consider several state-based continuous robotic control task distributions simulated via MuJoCo [48]. Following previous work [27, 29], we select the following task distributions: (1) Half-Cheetah-Fwd-Back, consisting of two opposite tasks; (2) Half-Cheetah-Vel and Humanoid-Direc-2D, each comprising 100 training tasks and 30 test tasks. More details of these task distributions are provided in Appendix D.
Baselines.
In DMC, we choose two model-free methods using Markovian policies $\Pi_{1}$: CURL [49] and SAC+AE [46]. In addition, we include two classic world models, PlaNet [4] and Dreamer [5], and two advanced contrastive learning-based methods, TACO [50] and TD-MPC [51], which utilize historical state-actions in policies and thus belong to $\Pi_{2}$. Moreover, we evaluate a SOTA model-based meta RL method belonging to $\Pi_{3}$: MAMBA [7]. For state-based control, besides PlaNet and Dreamer, we include meta-RL baselines such as MAML [27], RL2 [28], and VariBAD [31], including zero-shot and few-shot evaluation as references.
Metrics.
For the task combinations, we evaluate the average return of all tasks to verify TDR. For the task generalization setting, agents are trained on the training tasks and evaluated on the test tasks to assess their generalization performance. All experiments are repeated with 5 different random seeds, and results are reported as the mean $\pm$ std to mitigate the effects of randomness following previous works [5, 7].